-1/8(x-27)+1/3(x+3)=x-59 how to solve
1
Expert's answer
2020-09-15T15:26:09-0400

"-\displaystyle\frac{1}{8} (x - 27) + \frac{1}{3}(x + 3) = x - 59 \\

\textsf{Open the brackets and simplify}\\

\frac{-x}{8} + \frac{27}{8} + \frac{x}{3} + 1 = x - 59 \\

\frac{-x}{8} + \frac{x}{3} + \frac{35}{8} = x - 59 \\

\frac{-3x}{24} + \frac{8x}{24} + \frac{35}{8} = x - 59 \\

\frac{5x}{24} + \frac{35}{8} = x - 59 \\

\frac{5x}{24} - x = -59 - \frac{35}{8} \\

\frac{5x - 24x}{24} = \frac{-59 \times 8 - 35}{8} \\
\frac{19x}{24} = \frac{507}{8} \\
x = \frac{507}{8} \times \frac{24}{19} = \frac{1521}{19} \approx 80.05263"